# Microscopic parameters of a type-II superconductor measured by small-angle neutron scattering

### v1

@@ Abstract
A necessary condition for understanding and predicting the properties of any material is knowledge of microscopic parameters which control these properties. In superconductors these parameters are the radius of the orbital motion of electrons bound in Cooper pairs $R_{0}$ and the radius of the field-induced currents $r_{i}$ caused by precession of the pairs; one more parameter, associated with $r_{i}$, is the number density of Cooper pairs $n_{cp}$. In this communication we report on the first measurements of these parameters in a type-II superconductor (niobium) by SANS (small-angle neutron scattering). Other approaches potentially applicable for measuring the microscopic parameters are considered.
@@
Figure 1: (a) Equilibrium structure of the flux lines (FLs) in type-II superconductors in the mixed state. (b) Zoomed area with a FL passing through the network of micro-whirls; the FL takes the space originally occupied by the micro-whirl. $R_{\perp}$ is the rms radius of cylindrical volumes filled with CPs of different orientations; $r_{i}$ is the rms radius of micro-whirls equal to the radius of the field-induced currents in CPs and, respectively, to the radius of the FLs; $d$ is a parameter of the FL lattice (distance between neighboring scattering planes). Arrows in (b) depict induced currents. H is the intensity of the magnetic field directed into the page.
@@
Figure 2: (a) magnetization curves measured with a dcdc SQUID magnetometer on the twin’s sample used in this work. (b) the sample/field configuration in the SANS experiment. $H_{0}$ is the applied field directed parallel to the sample plane.
@@
Therefore, being in the field H (unlike the induction B of a magnetic field, its intensity or strength H is never equal to zero) CPs precess with Larmor frequency, forming field-induced circular current loops residing in transverse to H planes. A squared rms radius of the loops is $r_{i}^{2}=m_{cp}c^{2}/\pi n_{cp}(2e)^{2}$, where $m_{cp}$ and $n_{cp}$ are mass and number density of CPs, and $e$ is the unit charge; due to its origin (Larmor precession), $r_{i}$ does not depend on the field, as it also follows from the experiment muSR-2025 . Each loop, i.e., the current induced in each CP, has a negative magnetic moment, which leads to the perfect diamagnetism of the S phase in pinning-free ellipsoidal samples.
To minimize the free energy of the ensemble of CPs, the induced currents line up in an ideally ordered 2D hexagonal lattice of micro-whirls oriented parallel to H. Hence, the entropy of the pairs’ ensemble is zero, which is confirmed by the absence of thermoelectric effects in superconductors, first discovered by Meissner Meissner27 . In turn, zero entropy implies zero temperature of the ordered CPs, one of the consequences of which is zero resistivity of the S phase at under-critical total current.
The micro-whirls represent long solenoids of radius $r_{i}$ very tightly ”wound“ around the $H$-lines (a longitudinal distance between the neighboring current loops equals $2e^{2}/mc^{2}\approx 5\cdot 10^{-13}$ cm, about tripled diameter of the proton). Consequently, micro-whirls do not interact, which complies with the requirement of thermodynamics. On the other hand, in the case of inhomogeneous magnetization, the FLs pass through the sample via normally conducting holes in the network of superconducting micro-whirls. Herewith geometry of the FLs in type-II superconductors reproduces the geometry of the original micro-whirls’ structure. The cross-section of the equilibrium structure of the FLs and a zoomed area with one FL are depicted in Fig 1.
The radius of the orbital motion of paired electrons $R_{0}$, more specifically, its rms projection onto the transverse to H plane $R_{\perp}=R_{0}\sqrt{2/3}$, and the rms radius of the field-induced currents $r_{i}$ are key microscopic parameters controlling properties of superconductors in a state of thermodynamic equilibrium. In particular, if $r_{i}>R_{\perp}$, then the superconductor is of type II and vice versa Note on IMS . At the same time, the ratio $r_{i}/R_{\perp}\equiv\aleph$ is a constant of material depending on its purity. One can show that in type-II materials $\aleph=H_{c2}/H_{c}$, where $H_{c}$ is the thermodynamic critical field. A detailed description of the outlined micro-whirls model (MWM) with all necessary references is available in VK_book , a shorter version can be found in Encyclopedia .
@@
Sample. The sample was a one-side polished 1 mm thick single crystal Nb disc 19 mm in diameter fabricated in the Surface Preparation Laboratory, the Netherlands. It was annealed at 800 ℃  for 3 hours and electropolished. The applied magnetic field was set parallel to the stationary sample and perpendicular to the neutron beam. The sample was cooled in zero field, then the field was varied in one direction. The sample/field configuration and magnetization curves measured on a 7-mm in diameter twin’s sample are shown in Fig. 2.
@@
In the mixed state the situation was more favorable. As seen from Fig. 3, we observed three kinds of the diffraction patterns, with maxima of the first and second order, one maximum and no maximum. The exposition time per data-point was 0.5 hour. Experimental points for the intensity vs the scattering vector (log $I$ vs log $Q$) in vicinity of each maximum were approximated with the Gaussian curve from which the values of scattering vectors at the maxima $Q_{m1}$ and $Q_{m2}$, and one-sigma error bars were evaluated. Positions of $Q_{m1}$ and $Q_{m2}$ vs the applied field $H_{0}$, and calculated from them the FL lattice parameter $d$ are shown in Fig. 4.
After passing $H_{c1}$ from below the FLs start to fill the sample bulk, implying that more and more micro-whirls are replaced by FLs, and the lattice parameter $d$ decreases accordingly. As follows from Fig. 4b, the process is not linear: at first it resembles an avalanche, then continues with a gradually decreasing pace. This is consistent with the shape of the magnetization curves in Fig. 2a and with results of other authors, including those first reported by Shubnikov at al. Shubnikov-37_typeII ; the field dependence of the diffraction pattern (Fig. 4a) and the dynamics of observed variation of the lattice parameter (Fig. 4b) are consistent with SANS and STM data reported by Christen et al. Christen and Hess et al. Hess , respectively.
@@
Next, having $r_{i}$ one can compute the number density of Cooper pairs $n_{cp}=m_{cp}c^{2}/\pi r_{i}^{2}(2e)^{2}=mc^{2}/2\pi r_{i}^{2}e^{2}$, where $m$ is the free electron mass. This yields $n_{cp}=(3.3\pm 0.8)\cdot 10^{22}$ cm-3. Neglecting thermal expansion of Nb, number density of free (non-paired) electrons in this metal at zero temperature is $n_{e}=5.5\cdot 10^{22}$ cm-3 Grigoriev . Hence, $n_{cp}/n_{e}=0.6\pm 0.15=60\pm 15$ %. Thus, our data furnish a confirmation of the hypothesis of Gorter and Casimir of 1934 that at zero temperature all conduction electrons are in a condensed, i.e. paired, state Gorter_Casimir . Note that if, as one might assume, $r_{i}$ is equal to the London penetration depth, then $n_{cp}$ would be four times smaller.
Finally, the radius of the orbital motion of the paired electrons $R_{0}$ can be calculated from the definition of $\aleph$ and its relationship with $H_{c}$ and $H_{c2}$ for type-II superconductors. From this follows that $R_{0}=\sqrt{3/2}\,r_{i}/\aleph=\sqrt{3/2}\,r_{i}{H_{c}}/{H_{c2}}$. Hence, $R_{0}$ in Nb at 3.5 K equals $22\pm 5$ nm. It is worth noting that $R_{0}$ is the most “hidden” microscopic parameter and, accordingly, the least accessible for direct measurements (if such measurements are possible at all). On the other hand, according to the MWM, $R_{0}$ is the most fundamental parameter, since the other two ($r_{i}$ and $n_{cp}$) depend on the material purity.
Summary and outlook. We demonstrated an approach allowing to determine microscopic parameters of type-II superconductors from SANS patterns. The main advantage of this approach is that data obtained are conditioned by the properties of the sample bulk Sebastian . Accordingly, the data retrieved from SANS are indifferent to the state of the sample surface and therefore allow an objective assessment of the error.
The parameters reported in this paper are the radius of the field-induced microscopic currents in Cooper pairs, related to it the number density of the pairs, and the most “sacred” parameter, the radius of the orbital motion of the paired elections, i.e., the radius of Cooper pairs. All these values were acquired for the first time. The accuracy of the obtained quantities is within 25%; it can be improved by enlarging the exposure time.
@@
Alternatively, the microscopic parameters can be determined in a lesser time-consuming way from ESR and NMR measurements with samples in the Meissner state. To date, none of these resonances have been studied at such conditions. As expected, ESR spectra may have two resonances associated with precession of the bound and free electrons with $g$-factors 1 and 2, respectively. A primary parameter which can be extracted from these spectra is the number density of Cooper pairs $n_{cp}$. From there one can compute $r_{i}$ and, for type-II materials, $R_{0}$.
To find $R_{0}$ in type-I superconductors, one can use the approach applied by Shubnikov with coworkers Shubnikov-37_typeII and Pippard Pippard-53 . It consists in transforming a type-I superconductor into a type-II one by admixing an original type-I metal with a small amount ($\sim$ 0.1-1 wt.%) of another metal so that the condensation energy $E_{c}(=-\int\textbf{M}\cdot d\textbf{H})$ remains unchanged.
@@
Data  availability.   The  raw  SANS  data  are  available in  the  ISIS  repository  at  https://doi.org/10.5286/ISIS. E.RB2220299-4.
@@
  4. (4) V. Kozhevnikov, A.-M. Valente-Feliciano, T. Prokscha, P. J. Baker, A. Polyanskii, C. Van Haesendonck, in 16th Intentional Conference on Muon Spin Rotation, Relaxation and Resonance, St. John’s NL 2025; 30th International Conference on Low Temperature Physics, Bilbao, Spain, 2025.
@@
  6. (6) There is a special case when $r_{i}\approx R_{\perp}$ or $\aleph\approx 1$. Samples in this case exhibit simultaneously properties of type-I and type-II superconductors. Their magnetic structure consists of macroscopic (usually disordered) normal and superconducting domains, herewith the latter contain ordered FLs. The state of such samples is often referred to as an intermediate mixed state, which has been discussed in the literature since the 1970s; many of these discussions concern the properties of high-purity niobium samples (see, e.g., Brandt_2011 ; Sebastian_2022 ). A detailed study of niobium vs impurity content was reported in Prozorov_2024 .

### v2

@@ Abstract
A necessary condition for understanding and predicting the properties of any material is knowledge of microscopic parameters which control these properties in a state of thermodynamic equilibrium. One can show (see, e.g., Ref. VK_book ), that in superconductors these parameters are the radius of the orbital motion of electrons bound in Cooper pairs $R_{0}$ and the radius of the field-induced currents $r_{i}$ caused by precession of the pairs; one more parameter, associated with $r_{i}$, is the number density of Cooper pairs $n_{cp}$. In this communication we report on the first measurements of these parameters in a type-II superconductor (niobium) by SANS (small-angle neutron scattering). Other approaches to measuring the microscopic parameters are considered. Our work suggests novel avenues for studying superconductivity, important for disclosing its mechanisms in superconductors of all kinds.
Theoretical introduction.
@@
Figure 1: (a) Equilibrium structure of the flux lines (FLs) in type-II superconductors in the mixed state. (b) Zoomed area with a FL passing through the network of micro-whirls; the FL takes the space originally occupied by the micro-whirl. $R_{\perp}$ is the rms radius of cylindrical volumes filled with CPs of different orientations; $r_{i}$ is the rms radius of micro-whirls equal to the radius of the field-induced currents in CPs and, respectively, to the radius of the FLs; $d$ is a parameter of the FL lattice (distance between neighboring scattering planes). Arrows in (b) depict induced currents. H is the intensity of the magnetic field directed into the page. In this experiment $\textbf{H}=\textbf{H}_{0}$, the applied field.
Therefore, being in the field H (unlike the induction B of a magnetic field, its intensity or strength H is never equal to zero Maxwell ) CPs precess with Larmor frequency, thus forming field-induced circular current loops residing in transverse to H planes. Note that Larmor precession is the only type of motion of bound electrons that can meet the time-reversal symmetry mandatory for any equilibrium statistical system (the Second law of thermodynamics Planck ). At the same time the necessary condition for achieving thermodynamic equilibrium in magnetizing media is homogeneity of the field intensity H, which means that samples intended for studying equilibrium properties in magnetic field should have an ellipsoidal shape Maxwell .
A squared rms radius of the loops is $r_{i}^{2}=m_{cp}c^{2}/\pi n_{cp}(2e)^{2}$, where $m_{cp}$ and $n_{cp}$ are mass and number density of CPs, and $e$ is the unit charge ri vs lambda . Due to its origin (Larmor precession), $r_{i}$ does not depend on the field, as it also follows from the experiment VK_JSNM ; muSR-2025 . Each loop, i.e., the current induced in each CP, has a negative magnetic moment. This leads to ideal diamagnetism of the S-phase in thermodynamically equilibrium samples, as was discovered independently by Meissner and Ochsenfeld Meissner and Ochsenfeld , and by Ryabinin and Shubnikov Shubnikov-Rjabinin .
Figure 2: (a) magnetization curves measured by a dcdc SQUID magnetometer on a twin’s sample used in this work (i.e., fabricated from the same Nb rod and subjected to the same treatment). The sample was cooled at zero field. At an ascending applied field, that is in the zero-field cooled (ZFC) mode, this sample is in thermodynamic equilibrium, while at a descending field, or in the FC mode, it is away from equilibrium. The SANS experiment was carried out on the ZFC sample.
To minimize the free energy of the ensemble of CPs, the induced currents line up in an ideally ordered 2D hexagonal lattice of micro-whirls oriented parallel to H. Hence, the entropy of the pairs’ ensemble is zero, which is confirmed by the absence of thermoelectric effects in superconductors, first revealed by Meissner Meissner27 , as well as by the absence of temperature dependence of magnetization in diamagnets. In turn, zero entropy implies zero temperature of the ensemble of ordered pairs, one of the consequences of which is zero resistivity of the S phase at subcritical total current zero Tcp .
The micro-whirls represent long solenoids of radius $r_{i}$ very tightly ”wound“ around the H-lines (a longitudinal distance between the neighboring current loops equals $2e^{2}/mc^{2}\approx 5\cdot 10^{-13}$ cm, about tripled diameter of the proton). Consequently, micro-whirls do not interact with each other, which complies with requirement of thermodynamics. On the other hand, in the case of inhomogeneous magnetization, the FLs pass through the sample via normally conducting holes in the network of superconducting micro-whirls. Herewith geometry of the FLs in type-II superconductors reproduces the geometry of the original micro-whirls’ structure, which also follows from thermodynamics. The cross-section of the equilibrium structure of the FLs and a zoomed area with one FL are depicted in Fig 1. To appreciate the degree of FL ordering, we note that number density of FLs is of the order of $10^{8}$ mm-2.
The radius of the orbital motion of paired electrons $R_{0}$, more specifically, its rms projection onto the transverse to H plane $R_{\perp}=R_{0}\sqrt{2/3}$, and the rms radius of the field-induced currents $r_{i}$ are key microscopic parameters controlling properties of superconductors in a state of thermodynamic equilibrium. In particular, if $r_{i}>R_{\perp}$, then the superconductor is of type II and vice versa Note on IMS . At the same time, the ratio $r_{i}/R_{\perp}\equiv\aleph$ is a constant of material depending on its purity. One can show that in type-II materials $\aleph=H_{c2}/H_{c}$, where $H_{c}$ is the thermodynamic critical field.
The outlined micro-whirls model (MWM) is presented in detail in VK_book ; its shortened version is available in Encyclopedia . The original account of the MWM and experimental data that triggered its development are published in VK_JSNM . The MWM is based on the Bohr–Sommerfeld quantization condition applied to Cooper pairs; it is free of postulates and adjustable parameters. The MWM describes all known properties of superconductors and predicts some not yet known effects. The idea of micro-whirls in superconductors was first put forward by Hall Hall .
Figure 3: The sample/field configuration in the SANS experiment. $\textbf{H}_{0}$ is the applied field directed parallel to the sample plane.
@@
Experimental
@@
The sample was a one-side polished 1 mm thick single crystal Nb disc 19 mm in diameter fabricated in the Surface preparation laboratory, the Netherlands. It was annealed at 800 ℃  for 3 hours and electropolished. The applied magnetic field was parallel to the stationary sample and perpendicular to the neutron beam. The sample was cooled in a zero field, then the field was turned on and varied in one direction, starting from zero. At such conditions the flux trapped in the sample is negligible and therefore, as was first shown by Shubnikov with coworkers Shubnikov-37_typeII ; Shubnikov-Rjabinin , the sample of our geometry is in thermodynamic equilibrium even if its purity is not very high, as is the case in never perfectly clean type-II superconductors.
The magnetization curves measured on a 7-mm in diameter twin’s sample and sample/field configuration are shown in Figs. 2 and 3, respectively. As can be seen from Fig. 2, the magnetization data obtained in the ZFC mode are reproducible; in particular, it is easy to recognize that in the linear section of the magnetization curves the sample is in the Meissner state. This confirms the thermodynamic equilibrium of our sample in ZFC mode. Contrarily, in FC mode, i.e, at the descending field, the sample is out of equilibrium. It is worth noting that most of the previous SANS experiments have been performed with FC samples of transverse geometry Sebastian , where thermodynamic equilibrium is practically unattainable.
@@
In the mixed state the situation was more favorable. As seen from Fig. 4, we observed three kinds of the diffraction patterns, with maxima of the first and second order, one maximum and no maximum. The exposition time per data-point was half an hour. Experimental points for the intensity vs the scattering vector (log $I$ vs log $Q$) in vicinity of each maximum were approximated with the Gaussian curve from which the values of scattering vectors at the maxima $Q_{m1}$ and $Q_{m2}$, and one-sigma error bars were evaluated. Positions of $Q_{m1}$ and $Q_{m2}$ vs the applied field $H_{0}$, and calculated from them the FL lattice parameter $d$ are shown in Fig. 5.
After passing $H_{c1}$ from below the FLs start to fill the sample bulk, implying that more and more micro-whirls are replaced by FLs, and the lattice parameter $d$ decreases accordingly. As follows from Fig. 5b, the process is not linear: at first it resembles an avalanche, then continues with a gradually decreasing pace. This is consistent with the shape of the magnetization curves in Fig. 2 and with results of other authors, including those first reported by Shubnikov at al. Shubnikov-37_typeII ; the field dependence of the diffraction pattern (Fig. 5a) and the dynamics of observed variation of the lattice parameter (Fig. 5b) are consistent with SANS and STM data reported by Christen et al. Christen and Hess et al. Hess , respectively.
@@
Next, having $r_{i}$ one can compute the number density of Cooper pairs $n_{cp}=m_{cp}c^{2}/\pi r_{i}^{2}(2e)^{2}=mc^{2}/2\pi r_{i}^{2}e^{2}$, where $m$ is the free electron mass. This yields $n_{cp}=(3.3\pm 0.8)\cdot 10^{22}$ cm-3. Neglecting thermal expansion of Nb, number density of free (non-paired) electrons in this metal at zero temperature is $n_{e}=5.5\cdot 10^{22}$ cm-3 Grigoriev . Hence, $n_{cp}/n_{e}=0.6\pm 0.15=60\pm 15$ %. Thus, our data furnish a confirmation of the hypothesis of Gorter and Casimir of 1934 that at zero temperature all conduction electrons are in a condensed, i.e. paired, state Gorter_Casimir . Note that if, as someone might assume, $r_{i}$ is equal to the London penetration depth, then $n_{cp}$ would be four times smaller, which is quite unnatural.
Finally, the radius of the orbital motion of the paired electrons $R_{0}$ can be computed from the above definition of $\aleph$ and its relationship with $H_{c}$ and $H_{c2}$ for type-II superconductors. From this follows that $R_{0}=\sqrt{3/2}\,r_{i}/\aleph=\sqrt{3/2}\,r_{i}{H_{c}}/{H_{c2}}$. Hence, $R_{0}$ in Nb at 3.5 K equals $22\pm 5$ nm. It is worth noting that $R_{0}$ is the most “hidden” microscopic parameter and, accordingly, the least accessible for direct measurements (if such measurements are possible at all). On the other hand, according to the MWM, $R_{0}$ is the most fundamental parameter, since the other two ($r_{i}$ and $n_{cp}$) depend on the sample purity.
Summary and outlook.
We demonstrated an approach allowing to determine equilibrium microscopic parameters of type-II superconductors from SANS patterns. The main advantage of this approach is that data obtained are conditioned by the properties of the sample bulk. Accordingly, the data retrieved from SANS are indifferent to the state of the sample surface and therefore allow an objective assessment of the error.
The parameters reported are the radius of the field-induced microscopic currents in Cooper pairs, related to it the number density of the pairs, and the most “sacred” parameter, the radius of the orbital motion of the paired elections, i.e., the radius of Cooper pairs. All these values were acquired for the first time. The accuracy of the obtained quantities is within 25%; it can be improved by enlarging the exposure time.
@@
Alternatively, the microscopic parameters can be determined in a lesser time-consuming way from ESR and NMR measurements with homogeneous samples, i.e., with the samples in the Meissner state. To date, none of these resonances have been studied at such conditions. As expected, ESR spectra may have two resonances associated with precession of the bound and free electrons with $g$-factors 1 and 2, respectively. A primary parameter which can be extracted from these spectra is the number density of Cooper pairs $n_{cp}$. From there one can compute $r_{i}$ and, for type-II materials, $R_{0}$. The sample size in such cases may be significantly smaller than that required for SANS (but in any case the shape should be as ellipsoidal as possible).
To find $R_{0}$ in type-I superconductors, one can use the approach applied by Shubnikov et al. Shubnikov-37_typeII and Pippard Pippard-53 . It consists in transforming a type-I superconductor into a type-II one by admixing an original type-I metal with a small amount ($\sim$ 0.1-1 wt.%) of another metal so that the condensation energy $E_{c}(=-\int\textbf{M}\cdot d\textbf{H})$ remains unchanged.
@@
Data availability
The original data are available in the ISIS repository at https://doi.org/10.5286/ISIS.E.RB2220299-4.
@@
  4. (4) J. C. Maxwell, A Treatise on Electricity and Magnetism, v. II (Clarendon Press, Oxford, 1873).
  5. (5) Max Planck, Treatise on Thermodynamics, Third ed. (Dover Publication, N. Y. 1969).
  6. (6) Note that $r_{i}$ is numerically equal to twice the London penetration depth, but the physical meaning of $r_{i}$ is unrelated to the connstat of exponential decay of the magnetic field in the London theory.
  7. (7) V. Kozhevnikov, J Supercond Nov Magn 34, 1979 (2021).
  8. (8) V. Kozhevnikov, A.-M. Valente-Feliciano, T. Prokscha, P. J. Baker, A. Polyanskii, C. Van Haesendonck, 16th Intentional Conference on Muon Spin Rotation, Relaxation and Resonance, St. John’s NL 2025; 30th International Conference on Low Temperature Physics, Bilbao, Spain, 2025.
  9. (9) W. Meissner and R. Ochsenfeld, Naturwissenschaften 21, 787 (1933).
  10. (10) G. N. Rjabinin and L. W. Shubnikov, Nature 134, 286 (1934).
@@
  12. (12) Zero temperature of the ensemble of Cooper pairs means that the ordered micro-whirls do not interact with the environment. Hence, in the total current, Cooper pairs move all together at the drift velocity relative to the ionic lattice experiencing no resistance.
  13. (13) There is a special case when $r_{i}\approx R_{\perp}$ or $\aleph\approx 1$. Samples in this case exhibit simultaneously properties of type-I and type-II superconductors. Their magnetic structure consists of macroscopic (usually disordered) normal and superconducting domains, herewith the latter contain ordered FLs. The state of such samples is often referred to as an intermediate mixed state (see, e.g., Brandt_2011 ; Sebastian_2022 ). A detailed study of niobium vs impurity content was reported in Prozorov_2024 .